OpenStudy (anonymous):

could someone help me with this problem (-12)2

OpenStudy (aaronq):

is it \[(-12)^2\]

OpenStudy (aaronq):

or (-12)*2

OpenStudy (anonymous):

yes it's the first one

OpenStudy (aaronq):

when the sign is inside the brackets you include it in the computation (-12)^2 = (-12)*(-12)= 144 when it's outside like -(12)^2 then you only apply the negative sign at the end, eg -(12)^2=-144

OpenStudy (anonymous):

okay thanks for your help

OpenStudy (aaronq):

no problem